An electron is moving with a velocity V parallel to an infinitely long wire with uniform charge density (λ) and current I. what is V so that electron continues in a straight line? ( Please be really detailed about the methods and formulas used as I am so lost)
Answer #1

magnetic field due to current carrying wire

B = (uo/ 4pi) 2 I / a

electric field due to charge linear conductor

E = (1/ 4 pi eo) / a

for no deflection

E = B v

v = E/B

v = (1/4 pi eo) / ( (uo / 4pi) * 2I)

=======

Comment before rate in case any doubt, will reply for sure.. goodluck